is this the same as the section as
Applications of Linear Systems?
Ex: A motel clerk counts his $1 and $10 bills at the end of the day. He finds that he has a total of 74 bills having a combined monetary value of $326. Find the number of bills of each denomination that he has.
let x= the number of $1 bills the clerk has, and y= the number of $10 bills he has
He has 74 bills altogether, so one equation is x+y=74. To find the second equation, the dollar amount in $1 bills plus the dollar amount in $10 bills is $326, so 1x+10y=326.

Answer by rfer(16322)

x+y=74 y=74-x
---------------
x+10y=326
x+10(74-x)=326
x+740-10x=326
-9x=-414
x=-414/-9
x=46
------------
y=74-46
y=28